OpenStudy (anonymous):

The finction A=283e^(0.028t) models the population of a particular city, in thousands, t years after 1998. When will the city reach 335 thousand?

OpenStudy (harsimran_hs4):

simple just need to solve 283e^(0.028t) > 335 so e^(0.028t) > 335/283 solve this equation... if problem persists please ask

OpenStudy (anonymous):

Why is it greater than rather than equals?

OpenStudy (harsimran_hs4):

oops sorry i just read the question wrong it`s equal instead of = (i tought cross 335k mark)

OpenStudy (harsimran_hs4):

* instead of >

OpenStudy (anonymous):

Okay! (: How would you keep solving it out with the e involved?

OpenStudy (harsimran_hs4):

to remove e take ln on both sides

OpenStudy (dumbcow):

take log of both sides \[\ln e^{.028t} = \ln \frac{335}{283}\] \[.028t = \ln \frac{335}{283}\]

OpenStudy (anonymous):

Aye, dumbcow, the ln wouldn't stay on the left side?

OpenStudy (harsimran_hs4):

so what it`s constant on right side....you can find the value

OpenStudy (anonymous):

Do you get 8 years?

OpenStudy (anonymous):

Nevermind, I messed up! It's 6 years. Thanks! (:

OpenStudy (harsimran_hs4):

yes approximately 6yrs :)